# RuleRAG: Rule-Guided Retrieval-Augmented Generation with Language Models for Question Answering

### v1

@@ 3.2 RuleRAG-FT
The overview of our proposed rule-guided retriever and generator fine-tuning in RuleRAG-FT are illustrated in Figure 2 (b) and (c), respectively. For rule-guided retriever fine-tuning (RGFT-retriever), we update the LM encoders in a contrastive learning objective (Chen et al., 2020) and train over supervised fine-tuning data $\mathcal{F}_{R}$ provided in our constructed benchmarks, where inputs are the queries plus rules and supervised labels are heuristic oracle documents. Compared with retrievers employed with simple retrieval principles, our fine-tuned retrievers can recall more relevant results, aligned with the preferences of the rules. For rule-guided generator fine-tuning (RGFT-generator), we adopt the supervised instruction-tuning objective (Iyer et al., 2023; Chung et al., 2024) while combining each query $q$ with two components: retrieved documents $\mathcal{D}_{q}$ from the retrieval phase and the same set of rules $\rm{R}_{q}$ consistent with the retrieval phase. The rules introduced in the RGFT-generator train LLMs on how to optimally attribute from the retrieved context into answers by following rules, making RuleRAG leverage the fine-tuned retrievers more rationally. Experiments show our proposed RGFT can further guarantee and boost the retrieval quality and answering accuracy of RuleRAG-FT than RuleRAG-ICL.
@@
Rule-guided retriever fine-tuning (RGFT-retriever). We utilize two main types of retrievers: sparse retrievers and dense retrievers. As the sparse retriever, we use Pyserini  [^1] https://github.com/castorini/pyserini to implement the standard training-free BM25 (Robertson & Zaragoza, 2009), which relies on word-level frequencies. As the dense retrievers, we adopt the dual-encoder based retriever architecture, such as DPR  [^2] https://github.com/facebookresearch/DPR and SimCSE  [^3] https://github.com/princeton-nlp/SimCSE. We freeze the document encoder and tune the query encoder for high retrieval efficiency (Lewis et al., 2020). Given a $(\left(q,r\right),\mathcal{D}_{o})$ pair in the fine-tuning data $\mathcal{F}_{R}$ where $\mathcal{D}_{o}$ serve as the oracle documents, each $d_{i}^{+}\in\mathcal{D}_{o}$ is a positive learning example while each in-batch $d_{j}^{-}\not\in\mathcal{D}_{o}$ is a negative example. We train the retrievers in an in-batch contrastive training fashion with the following loss function $\mathcal{L}^{r}_{q}$:
$\mathcal{L}^{r}_{q}=-\log\frac{\exp(s({d}_{i}^{+},q\circ r))}{\exp(s({d}_{i}^{+},q\circ r))+\sum_{d_{j}^{-}\epsilon\mathcal{B}/\mathcal{D}_{o}}\exp(s({d}_{j}^{-},q\circ r))},\vspace{-0.02cm}$ (3)
@@
where $\mathcal{B}$ represents the documents for all the queries in one training batch. $\mathcal{D}_{o}$ represents oracle documents for the query and $\mathcal{B}/\mathcal{D}_{o}$ represents its in-batch negative examples. Retrievers are fine-tuned over $\mathcal{F}_{R}$. The training goal of RGFT-retriever is to minimize the overall loss $\mathcal{L}={\textstyle\sum_{(\left(q,r\right),\mathcal{D}_{o})\in\mathcal{F}_{R}}}\mathcal{L}^{r}_{q}$.
@@
Rule-guided generator fine-tuning (RGFT-generator). From RuleRAG-ICL, we find LLMs have a certain in-context learning ability to understand the rules. For greater model efficiency and control of the output, we fine-tune our generators in RuleRAG-FT and further enhance the proficiency of LLMs to attribute accurate answers following the instruction prompt. Formally, the designed instruction contains three parts: the relevant facts $\mathcal{D}_{q}$ retrieved by retrievers fine-tuned above, the rules $\rm{R}_{q}$ guiding attributable retrieval logics and the original query $q$. The instruction prompt remains the same during the fine-tuning of generators and inferring of RuleRAG to keep a similar knowledge distribution.
@@
In practice, for open-source LLMs, we utilize the few-shot instruction fine-tuning strategy considering the following two aspects. First, our introduced rules reform the data-centric training to the alignment of task-centric abilities, i.e., it can be viewed as a reasoning task based on the guidance of rules (Zhou et al., 2023) and our training aim is to learn to use them. Secondly, tuning all the data in $\mathcal{F}_{G}$ is prohibitive in time. We randomly select a fixed number of samples from $\mathcal{F}_{G}$ to conduct few-shot tuning (2048 samples in our practice). Our experiments show the effectiveness and generalization of RuleRAG-FT although the samples can not cover all the rules. For closed-source LLMs, we perform 3-shot prompts as an empirical substitute of fine-tuning (Dai et al., 2023) due to the unavailable parameters. Specifically, we randomly select three $((q,\mathcal{D}_{q},\rm{R}_{q}),$ $a)$ pairs from $\mathcal{F}_{G}$ as fixed examples in the prompts, making up the in-context augmentation. The detailed prompts are in the Appendix A.9.
@@ 4 Experimental Settings
4.1 Setup of RuleRAG
For our proposed RuleRAG-ICL, in addition to adding rule guidance to both retrievers and generators (RG-retriever + RG-generator), we also add rule guidance only to the retrieval stage (RG-retriever + generator), trying to prove that introducing rules in two stages can both contribute to the performance. For our proposed RuleRAG-FT, the complete method involves retrievers and generators with RGFT. The ablation study shows both of them are individually beneficial to the results. To emphasize the contribution of rules, we introduce several variants of RuleRAG-FT. The SSFT in Table 2 represents the standard supervised fine-tuning following the vanilla manner, where the fine-tuning instruction consists only of the queries and retrieved documents without rules. Note that whether or not the inputs are added with rules during inference is consistent with how the models are fine-tuned during training.
4.2 Baselines
Given that LLMs have pre-trained with lots of world knowledge, we report the performance of directly using LLMs as answer predictors without retrieval (Standard Prompting in Table 2) for basic performance reference (Ouyang et al., 2024). Additionally, we compare RuleRAG with a wide range of baselines based on retrieval-augmented generation (RAG). We instantiate the standard RAG framework using off-the-shelf LLMs and retrievers with queries as input, standing for the widespread RAG methods (Standard RAG in Table 2 and  3). Chain-of-thought (CoT) methods, verify-and-edit (VE;  Zhao et al. (2023)) and chain-of-knowledge (CoK;  Li et al. (2024)) correct LLM outputs independently and sequentially respectively by leveraging external knowledge sources. Following their implementation, we initialize the knowledge sources as our corpus $\mathcal{D}$ for a fair comparison and use 3-shot CoT prompts.
4.3 Evaluation Metrics
5 Experimental Results
5.1 Main Results
For RuleRAG-ICL (RG-DPR + LLAMA2_7B), introducing rules in the retrieval stage alone enhances the recall performance of the retriever and further improves the answer accuracy of the original LLAMA2_7B. RuleRAG-ICL (RG-DPR + RG-LLAMA2_7B) consistently surpasses Standard RAG across various metrics (+9.3 in R@10, +5.9 in EM and +3.2 in T-F1 on average absolute performance over all five benchmarks), achieving the improved performance. This confirms the sub-optimal ability of the current RAG and the effectiveness of our proposed dual rule-guided retriever and generator. For RuleRAG-FT, our proposed RGFT can amazingly improve performance by a significant margin (+45.7 in R@10, +24.2 in EM and +15.3 in T-F1 compared to the best performance of RuleRAG-ICL). To further corroborate that these gains are due to the introduced rules, we first isolate the key component, rules, from fine-tuning data $\mathcal{F}_{R}$ for RGFT, to form the standard supervised fine-tuning (SSFT) (Rule Ablation in Table 2) and then isolate the impact of the fine-tuned generator from the fine-tuned retriever in RuleRAG-FT (RGFT Ablation in Table 2). RGFT Ablation shows both RGFT-DPR and RGFT-LLAMA2_7B are beneficial when used individually, implicitly suggesting that the two phases do not depend on each other. Moreover, Rule Ablation shows when we no longer leverage rules to explicitly inform the retrievers of the retrieval directions (SSFT-DPR) or how LLMs should correctly utilise the retrieved documents while fine-tuning (SSFT-LLAMA2_7B), our recall and generation performances show varying degrees of degradation compared to RuleRAG-FT. This further clarifies the great assistance of rules on our method’s ability to answer knowledge-intensive queries.
Table 2: Performance comparison of RuleRAG-ICL and RuleRAG-FT with their variants and baselines. RG-DPR and RG-LLAMA2_7B represent rule-guided DPR and rule-guided LLAMA2_7B in RuleRAG-ICL. RGFT represents rule-guided fine-tuning in RuleRAG-FT. SSFT represents standard supervised fine-tuning (Section 4.1). Standard Prompting does not have a retrieval stage, so there is no R@10. VE and CoK involve multiple search objects, which change several times, so the R@10 loses reference value. The best performance of RuleRAG-ICL and RuleRAG-FT are in bold.
@@ 5.1 Main Results
5.2 Results of More LLMs
@@
To test the generalizability to more generators in RuleRAG-ICL and RuleRAG-FT, we evaluate how different LLMs affect the performance in Table 3. We experiment with three more open-source LLMs: ChatGLM2_6B (Du et al., 2022), Mistral_7B_v0.2 (Jiang et al., 2023a), LLAMA2_13B (Touvron et al., 2023), and a closed-source LLM, GPT-3.5-Turbo  [^4] https://openai.com/index/gpt-3-5-Turbo-fine-tuning-and-api-updates/, which can be called through OpenAI API.
@@
First, consistent with the conclusions for the LLAMA2_7B in Table 2, the results in Table 3 show RuleRAG is effective under various kinds of LLMs. RuleRAG-ICL and RuleRAG-FT improve the overall performance of Standard RAG across all benchmarks and LLMs, demonstrating the validity and universality of rules. RuleRAG-FT consistently outperforms RuleRAG-ICL. Secondly, for LLAMA2 as generators, Standard RAG, RuleRAG-ICL and RuleRAG-FT with the 13B model always outperform their 7B counterparts, indicating that the introduced rules can provide better guidance when using larger models with the same LLM architecture. Thirdly, take LLAMA2_13B as an instance, the EM results of RuleRAG-ICL with GPT-3.5-Turbo are better than RuleRAG-ICL with LLAMA2_13B for the more massive model parameters, however, the EM results of RuleRAG-FT with LLAMA2_13B are better than RuleRAG-FT with GPT-3.5-Turbo in three of the five benchmarks. This phenomenon illustrates that RGFT is fairly effective and necessary for lightweight LLMs, making RuleRAG-FT much cheaper than off-the-shelf big LLMs for LLM deployment and application.

### v2

@@ 3.2 RuleRAG-FT
The overview of our proposed rule-guided retriever and generator fine-tuning is shown in Figure 2 (b) and (c). For rule-guided retriever fine-tuning (RGFT-retriever), we update the LM encoders in a contrastive learning objective (Chen et al., 2020) and train over supervised fine-tuning data provided in our constructed benchmarks , where inputs are the queries plus rules and supervised labels are heuristic oracle documents. For rule-guided generator fine-tuning (RGFT-generator), we adopt the supervised instruction-tuning objective (Iyer et al., 2023) while combining query $q$ with two components: retrieved documents $\mathcal{D}_{q}$ and the set of rules $\rm{R}_{q}$ consistent with the retrieval phase. The rules introduced in the RGFT-generator train LLMs to optimally reason from the retrieved context into answers via attributable rules, making RuleRAG leverage our fine-tuned retrievers more rationally.
@@
Rule-guided retriever fine-tuning (RGFT-retriever). We utilize two main types of retrievers: sparse and dense retrievers. As the sparse retriever, we use Pyserini to implement the standard training-free BM25 (Robertson and Zaragoza, 2009), which relies on word-level frequencies. As the dense retrievers, we adopt the dual-encoder based retriever architecture, such as DPR and SimCSE. We freeze the document encoder and tune the query encoder for high retrieval efficiency (Lewis et al., 2020). Given a $(\left(q,r\right),\mathcal{D}_{o})$ pair in the fine-tuning data, where $\mathcal{D}_{o}$ serve as the oracle documents, each $d_{i}^{+}\in\mathcal{D}_{o}$ is a positive learning example while each in-batch $d_{j}^{-}\not\in\mathcal{D}_{o}$ is a negative example. We train the retrievers in an in-batch contrastive training fashion with the following loss function $\mathcal{L}^{r}_{q}$:
$\mathcal{L}^{r}_{q}=-\log\frac{e^{s({d}_{i}^{+},q\circ r)}}{e^{s({d}_{i}^{+},q\circ r)}+\sum_{d_{j}^{-}\epsilon\mathcal{B}/\mathcal{D}_{o}}e^{s({d}_{j}^{-},q\circ r)}},\vspace{-0.02cm}$ (3)
@@
where $\mathcal{B}$ is the documents for all the queries in one training batch. $\mathcal{D}_{o}$ is oracle documents for the query and $\mathcal{B}/\mathcal{D}_{o}$ is its in-batch negative examples. The final training goal of RGFT-retriever is to minimize the overall loss $\mathcal{L}={\textstyle\sum_{(\left(q,r\right),\mathcal{D}_{o})\in\mathcal{F}_{R}}}\mathcal{L}^{r}_{q}$.
@@
Rule-guided generator fine-tuning (RGFT-generator). To obtain greater model efficiency, we fine-tune the generators in RuleRAG-FT, enhancing the proficiency to reason accurate answers following rules. Formally, the designed instruction contains three parts: the relevant facts $\mathcal{D}_{q}$ retrieved by retrievers fine-tuned above, the rules $\rm{R}_{q}$ guiding attributable retrieval logics and the original query $q$.
@@
In practice, for open-source LLMs, we utilize the few-shot instruction fine-tuning strategy considering the following two aspects. First, our introduced rules reform the data-centric training to the alignment of task-centric abilities, i.e., it can be viewed as a reasoning task based on the guidance of rules (Zhou et al., 2023) and our training aim is to learn to use them. Secondly, tuning all the data is prohibitive. We randomly select a fixed number of samples to conduct few-shot tuning (2048 samples in our practice). For closed-source LLMs, we perform 3-shot prompts as an empirical substitute of fine-tuning (Dai et al., 2023) due to the unavailable parameters. Specifically, we randomly select three $((q,\mathcal{D}_{q},\rm{R}_{q}),$ $a)$ pairs as fixed examples in the prompts, making up the in-context augmentation.
Benchmarks $|\left|\mathcal{R}\right|$$\left|\mathcal{D}\right|$ $\left|\mathcal{F}_{R}\right|$ $\left|\mathcal{F}_{G}\right|$ $\left|\rm{Q}\right|$ Source KG
Temporal RuleQA-I 557 77,508 6,594 7,440 1,559 ICEWS14
RuleQA-Y 99 243,633 28,153 22,765 1,864 YAGO
RuleQA-W 78 584,364 50,996 62,375 2,065 WIKI
Static RuleQA-F 367 49,088 8,082 9,645 1,233 FB15K-237
RuleQA-N 234 18,177 4,351 4,764 815 NELL-995
Table 1: The statistics of our constructed benchmarks RuleQA. $\left|\mathcal{R}\right|$, $\left|\mathcal{D}\right|$, $\left|\mathcal{F}_{R}\right|$, $\left|\mathcal{F}_{G}\right|$ and $\left|\rm{Q}\right|$ are the numbers of rules, documents in corpus, retriever fine-tuning pairs, generator fine-tuning pairs and test queries, respectively.
@@ 4 Experimental Settings
4.1 Benchmarks and Setup of RuleRAG
The construction process of our constructed five rule-aware benchmarks RuleQA are in Appendix A. The statistics of RuleQA are in Table 1. For RuleRAG-ICL, in addition to adding rule guidance to both retrievers and generators (RG-retriever + RG-generator), we also add rule guidance only to the retrieval stage (RG-retriever + generator), trying to prove that introducing rules in two stages can both contribute to the performance. For RuleRAG-FT, the complete method involves retrievers and generators with RGFT. The ablation study shows both of them are individually beneficial to the results. To emphasize the contribution of rules, we introduce several variants of RuleRAG-FT. The SSFT in Table 2 represents the standard supervised fine-tuning following the vanilla manner, where the fine-tuning instruction consists only of the queries and retrieved documents without rules. Whether the inputs are added with rules during inference is consistent with how the models are fine-tuned.
@@ 4.1 Benchmarks and Setup of RuleRAG
Table 2: Performance comparison of RuleRAG-ICL and RuleRAG-FT with their variants and baselines. RG-DPR and RG-LLAMA2_7B represent rule-guided DPR and rule-guided LLAMA2_7B in RuleRAG-ICL. RGFT represents rule-guided fine-tuning in RuleRAG-FT. SSFT represents standard supervised fine-tuning. Standard Prompting does not have a retrieval stage, VE and CoK involve multiple search objects, which change several times, so there is no R@10. The best performance of RuleRAG-ICL and RuleRAG-FT are in bold.
4.2 Baselines
Given that LLMs have lots of world knowledge, we report the performance of directly using LLMs as answer reasoners without retrieval (Standard Prompting in Table 2). Additionally, we compare RuleRAG with three baselines based on RAG. We instantiate the widespread RAG framework using off-the-shelf LLMs and retrievers with queries as input, standing for the standard RAG methods (Standard RAG in Table 2,  4 and  5). Chain-of-thought (CoT) methods, verify-and-edit (VE;  Zhao et al. (2023)) and chain-of-knowledge (CoK;  Li et al. (2024)) correct outputs independently and sequentially respectively by leveraging external knowledge sources. Following their implementation, we initialize the knowledge sources as our corpus $\mathcal{D}$ and use 3-shot CoT prompts. Moreover, since RuleRAG relies solely on rule guidance instead of other sophisticated techniques like reflection or interleave, we also focus on the performance comparison of RuleRAG with and without rules.
4.3 Evaluation Metrics
Figure 3: The Reacll@k and EM performance of RuleRAG-FT in RuleQA-I with different numbers of retrieved documents and under multiple circumstances: three settings in DPR (DPR, SSFT-DPR and RGFT-DPR), three settings in SimCSE (SimCSE, SSFT-SimCSE and RGFT-SimCSE) and one setting in BM25. Horizontal numbers over the pillars represent EM for bar charts and slanted numbers around the lines represent Recall@k for line charts.
5 Experimental Results
5.1 Main Results
For RuleRAG-ICL (RG-DPR + LLAMA2_7B), introducing rules in the retrieval stage alone enhances DPR recall performance and improves the answer accuracy of the original LLAMA2_7B. RuleRAG-ICL (RG-DPR + RG-LLAMA2_7B) consistently surpasses Standard RAG across various metrics (+9.3 in R@10, +5.9 in EM and +3.2 in T-F1 on average absolute performance over all five benchmarks), achieving the improved performance. This confirms the sub-optimal ability of the current RAG and the effectiveness of our proposed dual rule-guided retriever and generator. For RuleRAG-FT, our proposed RGFT can amazingly improve performance by a significant margin (+45.7 in R@10, +24.2 in EM and +15.3 in T-F1 compared to the best performance of RuleRAG-ICL).
To further corroborate that these gains are due to the introduced rules, we first isolate the key component, rules, from fine-tuning data for RGFT, to form the standard supervised fine-tuning (SSFT) (Rule Ablation in Table 2) and then isolate the impact of the fine-tuned generator from the fine-tuned retriever in RuleRAG-FT (RGFT Ablation in Table 2). RGFT Ablation shows both RGFT-DPR and RGFT-LLAMA2_7B are beneficial when used individually, implicitly suggesting that the two phases do not depend on each other. Moreover, Rule Ablation shows when we no longer leverage rules to explicitly inform the retrievers of the retrieval directions (SSFT-DPR) or how LLMs should correctly utilise the retrieved documents while fine-tuning (SSFT-LLAMA2_7B), our recall and generation performances show varying degrees of degradation compared to RuleRAG-FT. This further clarifies the great assistance of rules on the ability to answer knowledge-intensive queries.
5.2 Further Analysis on Retrievers
Retrievers in RuleRAG-FT
In Figure 3, we initialize RuleRAG-FT with more retrievers: dense retrievers DPR (Karpukhin et al., 2020), SimCSE (Gao et al., 2021) and training-free sparse retriever BM25 (Robertson and Zaragoza, 2009), and we use several retrieval configurations: retrievers without fine-tuning or with SSFT/RGFT while recalling different numbers of top-scored documents. Before fine-tuning, the Recall@k and EM performance of the three retrievers are comparable and each has their own performance characteristic, with no obvious advantages or disadvantages. For instance, DPR has the best Recall@10 and SimCSE has the best EM under top-10 documents before fine-tuning.
After fine-tuning, DPR consistently outperforms SimCSE and RGFT consistently outperforms SSFT. Specifically, under considering top-scored documents with the same k, for the two trainable dense retrievers, the RGFT version recalls more relevant information (R@k) than the SSFT version by a large margin, demonstrating the generality of the proposed RGFT across different retrievers. As a result, the EM scores of the generated answers are better when higher-quality documents from retrievers are provided. Moreover, when the retrievers and generators are applied with RGFT, RuleRAG-FT shows substantial performance gains, even with the retrieval number limited to top-1. For DPR and SimCSE, as we include more documents, the Recall@k and EM scores increasingly improve. This shows that leveraging rules to guide the retrieval and generation processes builds a bridge between queries and answers since rules provide retrieval directions and attributable mechanisms. For BM25, although Recall@k keeps increasing, EM experiences a drop, probably due to the retrieved noise.
One additional finding is that even though the difference in R@2 between the original DPR and SimCSE is not large (10.1% vs 10.2%), the EM of generated answers can differ significantly (12.4% vs 9.1%). The reason may be that DPR’s retrieved content includes not only the correct answers but also other helpful information. RGFT further widens the gap of Reacll@k between DPR and SimCSE.
Retrievers in RuleRAG-ICL
Contriever (Izacard et al., 2022) is a powerful retriever with strong unsupervised performance and can transfer well to new applications. Therefore, it has been widely used in RAG. In Table 4 in Appendix B, we note that Contriever without the guidance of rules can achieve relatively good recall and RG-Contriever makes further enhancements. Compared to Standard RAG, RuleRAG-ICL with RG-Contriever and RG-generators also obtain varying degrees of performance improvement under the three LLMs. These results confirm the outstanding ability of our proposed rule-guided method.
5.3 Further Analysis on Generators
More LLMs as Generators
@@ 5.3 Further Analysis on Generators
To test RuleRAG’s generalization to more generators, we evaluate the effect of different LLMs in Table 5 in Appendix C. We experiment with three more open-source LLMs: ChatGLM2_6B (Du et al., 2022), Mistral_7B_v0.2 (Jiang et al., 2023a), LLAMA2_13B (Touvron et al., 2023), and a closed-source LLM, GPT-3.5-Turbo.
@@
First, consistent with the conclusions for LLAMA2_7B in Table 2, Table 5 show RuleRAG is effective under various kinds of LLMs. RuleRAG-ICL and RuleRAG-FT improve the overall performance of Standard RAG across all benchmarks and LLMs, demonstrating the validity and universality of rules. RuleRAG-FT consistently outperforms RuleRAG-ICL. Secondly, for LLAMA2 as generators, Standard RAG, RuleRAG-ICL and RuleRAG-FT with the 13B model always outperform their 7B counterparts, indicating that the introduced rules can provide better guidance when using larger models with the same LLM architecture. Thirdly, the RuleRAG-ICL’s EM results of GPT-3.5-Turbo are better than LLAMA2_13B because of more massive model parameters, however, the RuleRAG-FT’s EM results of LLAMA2_13B are better than GPT-3.5-Turbo in three of the five benchmarks. This phenomenon illustrates that RGFT is fairly effective and necessary for lightweight LLMs to overcome big LLMs, making RuleRAG-FT much cheaper than off-the-shelf big LLMs for LLM deployment and application.
Figure 4: The EM variation of RuleRAG-FT produces different characteristics due to the varying difficulty of the rules in our constructed five RuleQA benchmarks.
Impact of RGFT Data Volume
In Figure 4, the x-axis is the ratio of the amount of fine-tuned data to the total amount of data in RGFT. The y-axis is the ratio of EM performance to the optimal one under DPR and LLAMA2_7B, with closer to 100% indicating stronger performance. The overall performance trend of the five benchmarks is that the larger the amount of fine-tuning data, the better the results. Yet, since the different properties of the rules in different benchmarks lead to different degrees of difficulty in learning, the growth of model performance under different benchmarks exhibits various characteristics.
Specifically, RuleQA-F has the most intuitive growth curve: It starts from low performance and slowly grows to the optimal performance, reflecting the increasing mastery of rules during the RGFT process of RuleRAG-FT. RuleQA-W yields relatively superior performance after just one-eighth of the total amount of the fine-tuning data and further improves subsequently. RuleQA-N achieves the best EM performance after three-eighths of the fine-tuning data and maintains flat. In contrast, the performance in RuleQA-Y fluctuates modestly at a very low level throughout the first half of the RGFT process (from one-eighth to five-eighths), and then sees a sudden surge in capability during the second half of the RGFT process (from six-eighths to the end). The EM performance in RuleQA-I fluctuates more dramatically: While realizing very large EM performance gains (ranking second in all the benchmarks), it undergoes several upward and downward drops before levelling off at the optimal performance. This suggests that RuleQA-I is the most challenging among our constructed five benchmarks. Moreover, from Table 2,  4 and  5, we find RuleRAG has the worst absolute performance in RuleQA-I compared to the other four benchmarks under the same LLMs, which also illustrates the challenge of our constructed RuleQA-I.
Figure 5: The EM of generalizing RuleRAG-FT from the source rule bank $\mathcal{R}_{i}$ to the target rule bank $\mathcal{R}_{j}$ , i.e., RuleRAG-FT is trained on $\mathcal{R}_{i}$ and tested on $\mathcal{R}_{j}$. The numbers in ($\mathcal{R}_{i}$, $\mathcal{R}_{j}$) represent the performance gains compared to the baseline Standard RAG tested on $\mathcal{R}_{j}$.
6 Rule Generalization
Generalization on RuleQA
RuleRAG-ICL is training-free, so we can attach arbitrary rules to the method’s input by in-context learning. Experimental results above naturally illustrate its instruction-following ability to many kinds of rules. In RGFT, the constructed fine-tuning data is limited anyway but rules are inexhaustible, so RuleRAG-FT cannot and should not see the full set of rules in RuleQA. Therefore, it is important to verify its ability to generalize to unseen rules.
RuleRAG-FT must capture transferable rule utilization capability, since RuleRAG-FT has no prior knowledge of the target rule bank and is forced to learn from the source rule bank. The results in Figure 5, where $\mathcal{R}_{i}\cap\mathcal{R}_{j}=\emptyset$ and $\left|\mathcal{R}_{i}\right|=\left|\mathcal{R}_{j}\right|(i,j\in\{1,2,3,4\})$, show that (1) The diagonal ($\mathcal{R}_{i},\mathcal{R}_{i}$) has the highest performance gains and there are slight differences between various rule banks; (2) The results on two sides of the diagonal fluctuations within reasonable ranges and all show stable improvements over Standard RAG. This implies that RuleRAG-FT can take advantage of the ability to leverage the learned underlying rule patterns rather than being limited to concrete rule instances.
Datasets ASQA PopQA HotpotQA NQ
Methods EM EM EM EM
Contriever + LLAMA2_7B
Standard RAG 8.6 14.3 4.4 7.6
RuleRAG-ICL 10.0 15.3 5.4 7.8
RuleRAG-FT 11.1 16.7 6.0 8.1
Contriever + LLAMA2_13B
Standard RAG 8.8 13.7 5.8 7.8
RuleRAG-ICL 10.4 17.3 6.8 8.3
RuleRAG-FT 11.9 18.8 8.2 8.9
CoK 27.9 11.6 36.8 31.4
CoK with Rules 40.0 16.2 38.6 35.4
Table 3: The results of RuleRAG and CoK on four common datasets adaptively equipped with rules in RuleQA.
Generalization on More Datasets
To test RuleRAG’s performance on retrieving and reasoning using rules in a wider range of scenarios, we conduct assessments in four datasets: ASQA (Stelmakh et al., 2022), PopQA (Mallen et al., 2023), HotpotQA (Yang et al., 2018) and Natural Questions (NQ) (Kwiatkowski et al., 2019). Table 3 shows RuleRAG’s results on them.
Even though these datasets were constructed without adapting rules, RuleRAG still achieves consistent performance gains with the help of the rules in our constructed RuleQA. Specifically, following the framework of RuleRAG, existing datasets can be adaptively equipped with rules in RuleQA by calculating the relevance between candidate rules and queries. If some rules are highly relevant, they are introduced, otherwise no rules are introduced. Table 3 also compares the performance changes of CoK without and with our introduced rules, indicating that the guidance of rules is beneficial to the retrieval and generation accuracy of existing models. These results further confirm the effectiveness of our proposed rule-guided retrieval and generation in RAG for more comprehensive QA applications.
7 Conclusion and Future Works
In this paper, we point out two high-level problems of current RAG and propose rule-guided retrieval-augmented generation (RuleRAG). RuleRAG-ICL intuitively shows RAG can directly benefit from prompting LLMs with rules by in-context learning. To further improve the QA performance, RuleRAG-FT retrofits retrievers to recall more supportive information by contrastive learning and updates generators through our designed RGFT. Experiments on our constructed five rule-aware QA benchmarks RuleQA show the strong performance of RuleRAG under multiple retrievers and generators and the generalization of rules. Furthermore, the comparison results with and without rules in RuleQA for RuleRAG and CoK on existing RAG datasets also attest to the effectiveness of rules in broader scenarios. In the future, we will explore how to adapt rules in more complex RAG frameworks and use custom rules for more QA tasks.
Limitations
Since existing RAG datasets do not have adapted rules, which have been widely used for knowledge-intensive reasoning tasks, we use mature KG rule mining algorithms to match rules for our constructed benchmarks RuleQA. Although the experiments on four existing RAG datasets, including ASQA, PopQA, HotpotQA and NQ, initially demonstrated that the guideline of rules in RuleQA can be generalized to them and yielded performance gains, the gains were limited because the rules were not customized for them. Therefore, we plan to match rules for more RAG datasets and validate the rules on more RAG models to demonstrate the generic usefulness, since all RAG-based methods involve the two basic processes of retrieval and generation.
Appendix A Newly Constructed Rule-aware QA Benchmarks
Many real-world scenarios, such as healthcare, law and finance, rely on expert experience and the expertise can be represented with symbolic rules. For example, "having certain symptoms corresponds to a certain disease, which in turn requires the use of certain medications", "certain behaviours violate certain laws, which in turn require certain penalties", and "when approving a loan, information such as the borrower’s debt-income ratio must be taken into account", and so on. In a broad sense, human common sense, expert experience, or regulations are rules and are ubiquitous in real life, but the common RAG datasets that are available today do not provide a corresponding rule base. We find that knowledge bases and their rule mining algorithms can provide high-quality rules and question-answer pairs, so we construct RuleQA based on knowledge bases and equip rules to RuleQA.
Rule bank $\mathcal{R}$. A huge amount of world knowledge, including static facts and temporal events, has been stored in static KGs and temporal KGs (Jiang et al., 2023b). If Event1 can lead to the happening of Event2, we believe that there is a logical correlation between them. In KGs, events are usually stored in the form of triples [Entity 1, $r_{1}$, Entity 2] , so we leverage a triple to stand for an event. In the static scenario, several different relations can be simultaneously established between two entities. In the temporal scenario, two entities can interact multiple times at different timestamps. Hence, if relation $r_{1}$ (rule body) can logically explain the occurrence of relation $r_{2}$ (rule head) between entities, we represent this relevance as rule $r$ in a natural language form: [Entity 1, $r_{1}$, Entity 2] leads to [Entity 1, $r_{2}$, Entity 2]. We leverage the classical rule mining algorithm AMIE3 (Lajus et al., 2020) for static KGs and TLogic (Liu et al., 2022) for temporal KGs. AMIE3 and TLogic are currently the most widely used rule mining methods, as well as one of the best static and temporal knowledge graph mining models, respectively. Therefore, the quality of the rules is guaranteed. They can intuitively give rules with confidence scores. The frequently co-occur relations form rules with high confidence (Liao et al., 2024) and we only transform these high-confidence rules to the above text string form, comprising our rule bank $\mathcal{R}$, which will be consistently leveraged in the training and inferring process of RuleRAG.
Benchmarks with temporal queries, named RuleQA-I, RuleQA-Y and RuleQA-W, are constructed based on three temporal KGs, ICEWS14 (García-Durán et al., 2018), YAGO (Mahdisoltani et al., 2013) and WIKI (Leblay and Chekol, 2018). Benchmarks with static queries, named RuleQA-F and RuleQA-N, are constructed based on two static KGs, FB15K-237 (Toutanova and Chen, 2015) and NELL-995 (Xiong et al., 2017).
Appendix B Further Analysis on Retrievers
Table 4 shows the performance of RuleRAG-ICL with Contriever and five LLMs.
Architecture RuleQA-I RuleQA-Y RuleQA-W RuleQA-F RuleQA-N
RuleRAG-ICL RG-Contriever Mistral_7B_v0.2 45.5 12.9 22.7 55.2 38.4 37.5 63.2 44.2 45.0 83.9 23.9 38.2 88.5 35.1 27.0
RG-Contriever RG-Mistral_7B_v0.2 45.5 15.4 24.5 55.2 40.8 39.8 63.2 46.3 45.8 83.9 26.1 39.8 88.5 39.8 31.9
Standard RAG Contriever LLAMA2_13B 41.2 22.1 39.5 52.7 40.8 44.2 62.2 49.2 52.4 80.6 42.4 51.4 87.6 50.2 57.4
RuleRAG-ICL RG-Contriever LLAMA2_13B 45.5 22.3 39.6 55.2 41.1 44.4 63.2 49.9 52.9 83.9 45.0 52.0 88.5 51.1 57.6
RG-Contriever RG-LLAMA2_13B 45.5 22.3 39.8 55.2 41.5 45.8 63.2 51.2 54.2 83.9 46.6 52.2 88.5 52.7 58.1
Table 4: The performance of RuleRAG-ICL with a powerful retriever, Contriever, under five LLMs.
Appendix C Further Analysis on LLMs
Table 5 is the performance of RuleRAG-ICL and RuleRAG-FT with four more LLMs as generators.